OpenStudy (anonymous):

-3 5/8 + 7 3/4 = plz help WRITE AS A MIXED NUMBER :) WILL GIVE FAN AND A MEDAL

OpenStudy (anonymous):

7 3/4 - 3 5/8= 31/4-29/8=62/8-29/8=33/8=4 1/8